# CoMatcher: Multi-View Collaborative Feature Matching

### v1

@@ Abstract
This paper proposes a multi-view collaborative matching strategy for reliable track construction in complex scenarios. We observe that the pairwise matching paradigms applied to image set matching often result in ambiguous estimation when the selected independent pairs exhibit significant occlusions or extreme viewpoint changes. This challenge primarily stems from the inherent uncertainty in interpreting intricate 3D structures based on limited two-view observations, as the 3D-to-2D projection leads to significant information loss. To address this, we introduce CoMatcher, a deep multi-view matcher to (i) leverage complementary context cues from different views to form a holistic 3D scene understanding and (ii) utilize cross-view projection consistency to infer a reliable global solution. Building on CoMatcher, we develop a groupwise framework that fully exploits cross-view relationships for large-scale matching tasks. Extensive experiments on various complex scenarios demonstrate the superiority of our method over the mainstream two-view matching paradigm.
@@ 3 Methodology
Formulation. Given a set of $N_{I}$ images, =$\mathcal{I}=\{I_{i}\mid i=1,\ldots,N_{I}\}$, of a scene, the objective of feature matching is to identify the 2D position coordinates $\mathcal{M}_{k}=\{\mathbf{p}^{I_{i}}\}$ in $\mathcal{I}$ that correspond to identical 3D scene points $\mathcal{X}=\{\mathbf{x}_{k}\mid k=1,\ldots,N_{X}\}$, where $\mathbf{p}^{I_{i}}\in\mathbb{R}^{2}$ are the pixel coordinates in $I_{i}$, and each $\mathcal{M}_{k}$ is called a track.
@@
Constructing complete tracks requires estimating correspondences between each image $I_{t}\subset\mathcal{I}$ and its co-visible image set $\mathcal{C}_{t}$. While the pairwise approach offers implementation flexibility for this problem, its reliance on two-view inputs in a single estimation often leads to uncertainty in complex scenes. In practice, images within $\mathcal{C}_{t}$ are often not independent—most are captured from adjacent positions but with differing viewing angles. This inspires us to group $\mathcal{C}_{t}$ based on image correlations, where each complementary group offers a mapping from multiple planar views to a 3D sub-scene. In the image matching problem, the correspondence between $I_{t}$ and a group must adhere to certain 2D-to-3D physical constraints, such as cross-view projection consistency. Rather than independently estimating and then integrating them with constraints, it is more efficient to unify this process and learn underlying priors directly from the data. Thus, we formulate CoMatcher to address the $1$-to-$N$ matching problem (see Fig. 2).
@@ 3.2 Multi-view feature interaction
Additionally, we also interact with all target view features after source cross-attention (Target Cross). For a point $t$ in view $I_{t}$ paired with source view $I_{i}$, we attend to the same point $t$ across the remaining $M-1$ target views. This is analogous to operations in some works that aggregate features along the temporal dimension [12, 26]. Unlike the approach for source views, we compute the attention scores solely based on feature similarity. Despite its low computational complexity, this module complements the Source Cross, mutually enhancing each other to help the network implicitly infer globally optimal correspondences.
@@ 4.3 Homography estimation
Tab. 1 shows that CoMatcher yields more accurate estimates than all sparse two-view matchers, highlighting the advantage of inference in multi-view feature space. However, despite End2End also jointly reasoning over multiple views like ours, its performance is notably inferior. This underscores the superiority of a $N$-to-$1$ architecture over $N$-to-$N$ for multi-view representation learning. When compared to LoFTR, we observe slightly lower performance under low thresholds. We attribute this mainly to the limitations of the feature extractor in keypoint non-repeatability and localization errors [29]. Additionally, DLT achieves accuracy close to RANSAC on most metrics, reflecting the high-quality correspondences from CoMatcher.
@@ 4.6 Ablation study
Table 4: Ablation study on synthetic homography dataset and HPatches. The “full" model is the default model.
@@ A Groupwise matching pipeline details
To reiterate, our method constructs a set of tracks $\{\mathcal{M}_{k}\mid k=1,\ldots,N_{X}\}$, where each track $\mathcal{M}_{k}=\{\mathbf{p}^{I_{i}}\}$ corresponds to one of the $N_{X}$ 3D points observed in an unconstrained image set =$\mathcal{I}=\{I_{i}\mid i=1,\ldots,N_{I}\}$. As illustrated in Fig. 6, the pipeline proceeds sequentially: it first extracts local features from each image, followed by identifying overlapping image pairs. Subsequently, the process involves grouping, connecting, and matching. Finally, all pairwise matches undergo robust geometric verification [17, 9] before being merged into global tracks. Detailed descriptions of the key components are provided in the following.
@@ C.1 Architecture
[…]
@@ D.1 Training details
Finetuning on MegaDepth. The model is fine-tuned of MegaDepth [28] with pseudo ground-truth camera poses and depth images. We sample 200 co-visible multi-view quadruplets per scene, randomly select one image as the target view for training. The sampling process follows image pair co-visibility scores, with a designed strategy to ensure balanced distribution across score intervals [0.1, 0.3], [0.3, 0.5], and [0.5, 0.9]. Images are resized such that their longer edge is 1024 pixels and zero-padded to a $1024\times 1024$ resolution. Correspondences with a reprojection error $\leq 3$ pixels and mutual nearest neighbors are labeled as inliers, while those with a reprojection error $>5$ pixels are labeled as outliers. Points without depth or without a correspondence having a sampson error $\leq 3$ pixels are also marked as outliers. We extract 2048 keypoints per image.

### v2

@@ Abstract
This paper proposes a multi-view collaborative matching strategy for reliable track construction in complex scenarios. We observe that the pairwise matching paradigms applied to image set matching often result in ambiguous estimation when the selected independent pairs exhibit significant occlusions or extreme viewpoint changes. This challenge primarily stems from the inherent uncertainty in interpreting intricate 3D structures based on limited two-view observations, as the 3D-to-2D projection leads to significant information loss. To address this, we introduce CoMatcher, a deep multi-view matcher to (i) leverage complementary context cues from different views to form a holistic 3D scene understanding and (ii) utilize cross-view projection consistency to infer a reliable global solution. Building on CoMatcher, we develop a groupwise framework that fully exploits cross-view relationships for large-scale matching tasks. Extensive experiments on various complex scenarios demonstrate the superiority of our method over the mainstream two-view matching paradigm. Code is available at https://github.com/EATMustard/CoMatcher.
@@ 3 Methodology
Formulation. Given a set of $N_{I}$ images, $\mathcal{I}=\{I_{i}\mid i=1,\ldots,N_{I}\}$, of a scene, the objective of feature matching is to identify the 2D position coordinates $\mathcal{M}_{k}=\{\mathbf{p}^{I_{i}}\}$ in $\mathcal{I}$ that correspond to identical 3D scene points $\mathcal{X}=\{\mathbf{x}_{k}\mid k=1,\ldots,N_{X}\}$, where $\mathbf{p}^{I_{i}}\in\mathbb{R}^{2}$ are the pixel coordinates in $I_{i}$, and each $\mathcal{M}_{k}$ is called a track.
@@
Constructing complete tracks requires estimating correspondences between each image $I_{t}\subset\mathcal{I}$ and its co-visible image set $\mathcal{C}_{t}$. While the pairwise approach offers implementation flexibility for this problem, its reliance on two-view inputs in a single estimation often leads to uncertainty in complex scenes. In practice, images within $\mathcal{C}_{t}$ are often not independent—most are captured from adjacent positions but with differing viewing angles. This inspires us to group $\mathcal{C}_{t}$ based on image correlations, where each complementary group offers a mapping from multiple planar views to a 3D sub-scene. In the image matching problem, the correspondence between $I_{t}$ and a group must adhere to certain 2D-to-3D physical constraints, such as cross-view projection consistency. Rather than independently estimating and then integrating them with constraints, it is more efficient to unify this process and learn underlying priors directly from the data. Thus, we formulate CoMatcher to address the 11-to-NN matching problem (see Fig. 2).
@@ 3.2 Multi-view feature interaction
Additionally, we also interact with all target view features after source cross-attention (Target Cross). For a point $t$ in view $I_{t}$ paired with source view $I_{i}$, we attend to the same point $t$ across the remaining M1M-1 target views. This is analogous to operations in some works that aggregate features along the temporal dimension [12, 26]. Unlike the approach for source views, we compute the attention scores solely based on feature similarity. Despite its low computational complexity, this module complements the Source Cross, mutually enhancing each other to help the network implicitly infer globally optimal correspondences.
@@ 4.3 Homography estimation
Tab. 1 shows that CoMatcher yields more accurate estimates than all sparse two-view matchers, highlighting the advantage of inference in multi-view feature space. However, despite End2End also jointly reasoning over multiple views like ours, its performance is notably inferior. This underscores the superiority of a NN-to-11 architecture over NN-to-NN for multi-view representation learning. When compared to LoFTR, we observe slightly lower performance under low thresholds. We attribute this mainly to the limitations of the feature extractor in keypoint non-repeatability and localization errors [29]. Additionally, DLT achieves accuracy close to RANSAC on most metrics, reflecting the high-quality correspondences from CoMatcher.
@@ 4.6 Ablation study
Table 4: Ablation study on synthetic homography dataset and HPatches. The “full” model is the default model.
@@ A Groupwise matching pipeline details
To reiterate, our method constructs a set of tracks $\{\mathcal{M}_{k}\mid k=1,\ldots,N_{X}\}$, where each track $\mathcal{M}_{k}=\{\mathbf{p}^{I_{i}}\}$ corresponds to one of the $N_{X}$ 3D points observed in an unconstrained image set $\mathcal{I}=\{I_{i}\mid i=1,\ldots,N_{I}\}$. As illustrated in Fig. 6, the pipeline proceeds sequentially: it first extracts local features from each image, followed by identifying overlapping image pairs. Subsequently, the process involves grouping, connecting, and matching. Finally, all pairwise matches undergo robust geometric verification [17, 9] before being merged into global tracks. Detailed descriptions of the key components are provided in the following.
@@ C.1 Architecture
Here, $\cdot\mid\cdot$] denotes concatenation, and $\operatorname{MLP}$ represents a multi-layer perceptron. The message vector $\mathbf{m}_{u}^{I_{i}\leftarrow\mathcal{W}}$ is computed through an attention mechanism [54], representing feature interactions between point $u$ and all points in $\mathcal{W}$. The update $\operatorname{MLP}$ consists of a single hidden layer with a dimension of $d_{h}=2d$, followed by a LayerNorm operation, a GeLU activation, and a linear projection from $(2d,d)$ with a bias term.
@@ D.1 Training details
Finetuning on MegaDepth. The model is fine-tuned of MegaDepth [28] with pseudo ground-truth camera poses and depth images. We sample 200 co-visible multi-view quadruplets per scene, randomly select one image as the target view for training. The sampling process follows image pair co-visibility scores, with a designed strategy to ensure balanced distribution across score intervals [0.1, 0.3], [0.3, 0.5], and [0.5, 0.9]. Images are resized such that their longer edge is 1024 pixels and zero-padded to a $1024\times 1024$ resolution. Correspondences with a reprojection error $\leq 3$ pixels and mutual nearest neighbors are labeled as inliers, while those with a reprojection error >5>5 pixels are labeled as outliers. Points without depth or without a correspondence having a sampson error $\leq 3$ pixels are also marked as outliers. We extract 2048 keypoints per image.